# Notions of rank and independence in countably categorical theories

### v1

@@ Abstract.
For an $\omega$-categorical theory $T$ and model $\mathcal{M}$ of $T$ we define a hierarchy of ranks, the $n$-ranks for $n<\omega$ which only care about imaginary elements “up to level nn”, where level $n$ contains every element of $M$ and every imaginary element that is an equivalence class of an $\emptyset$-definable equivalence relation on $n$-tuples of elements from $M$. Using the $n$-rank we define the notion of $n$-independence. For all $n<\omega$, the $n$-independence relation restricted to $M_{n}$ has all properties of an independence relation according to Kim and Pillay [15] with the possible exception of the symmetry property. We prove that, given any $n<\omega$, if $\mathcal{M}\models T$ and the algebraic closure in $\mathcal{M}^{\text{\rm eq}}$ restricted to imaginary elements “up to level nn” which have $n$-rank 1 (over some set of parameters) satisfies the exchange property, then $n$-independence is symmetric and hence an independence relation when restricted to $M_{n}$. Then we show that if $n$-independence is symmetric for all $n<\omega$, then $T$ is rosy. An application of this is that if $T$ has weak elimination of imaginaries and the algebraic closure in $\mathcal{M}$ restricted to elements of $M$ of 0-rank 1 (over some set of parameters from $M^{\text{\rm eq}}$) satisfies the exchange property, then $T$ is superrosy with finite $U^{\text{\rm\th }}$-rank.
@@ 1. Introduction
A variety of notions of rank and independence have played an important role in model theory at least since Morley’s influencial work on uncountably categorical theories [19] in the 1960ies. Such notions have been central for developing more or less general (meta) theories which divide complete first-order theories into various classes. Shelah [24] used them in his development of stability theory by which theories can be classified into $\omega$-stable, superstable, stable, or unstable. Later Shelah’s classification theory, and its notion of independence, was shown, by Kim and Pillay [15, 14], to make sense in a wider context and the classes of simple and supersimple first-order theories were introduced. Yet later, more general notions of independence, including thorn-independence, have been studied by Onshuus, Ealy and Adler and the classes of rosy and superrosy theories have been introduced [1, 7, 21]. The class of rosy theories is quite diverse and includes, for example, all simple theories (which exclude a linearly ordered universe) and all o-minimal theories (which assume a linearly ordered universe) [28]. It is also the largest class of theories for which there is an independence relation which satisfies certain basic and natural properties [7]. The work on NIP theories [25] also uses notions of rank and independence (and NIP theories and rosy theories partially overlap).
@@
In this study we consider (only) $\omega$-categorical theories $T$ and investigate a hierarchy of ranks, the $n$-ranks for $n<\omega$, where the $n$-rank is defined entirely in terms of the algebraic closure operator on $\mathcal{M}^{\mathrm{eq}}$ (where $\mathcal{M}$ is a model of the theory) restricted to the set $M_{n}$ containing every real element and every imaginary element that corresponds to an equivalence class of an $\emptyset$-definable equivalence relation on $M^{k}=\underbrace{M\times\ldots\times M}_{\text{$k$ times}}$ where $k\leq n$. With the $n$-rank we define a notion of $n$-independence. Without further assumptions than $\omega$-categoricity, $n$-independence restricted to $M_{n}$ has all the properties of an independence relation according to [15] (possibly) except for the symmetry property. The main technical contribution of this study is to isolate a property, parametrized by $n$, which implies (in fact, is equivalent to) that $n$-independence is symmetric. The property in question (Assumption 5.1) is the exchange property of algebraic closure restricted to elements of $M_{n}$ with $n$-rank 1 (over some $C\subseteq M^{\mathrm{eq}}$). More precisely, if, for any $C\subseteq M^{\mathrm{eq}}$, the algebraic closure using parameters from $C$ restricted to elements of $M_{n}$ with $n$-rank 1 over $C$ has the exchange property, then $n$-independence is symmetric, and hence it is an independence relation in the sense of [15] when restricted to $M_{n}$, which is stated by Theorem 5.15. Using this we can relatively easily show the statement of Theorem 6.10 that if, for all $n<\omega$, the algebraic closure restricted to elements of $M_{n}$ with $n$-rank 1 has the exchange property, then the theory is rosy, which is proved by showing that thorn-independence has local character (see [7, Theorem 3.7]). Thus, an $\omega$-categorical theory which is not rosy must, for some $n<\omega$, have a model $\mathcal{M}$ and $C\subseteq M^{\mathrm{eq}}$ such that algebraic closure using parameters from $C$ restricted to elements of $M_{n}$ with $n$-rank 1 over $C$ does not have the exchange property. This may be useful for finding a dividing line between $\omega$-categorical rosy theories and $\omega$-categorical non-rosy theories. Actually, I am not aware of any example of an $\omega$-categorical theory that is not rosy. Hence the results presented here may help to either find an $\omega$-categorical non-rosy theory, or to rule out the existence of such a theory.
@@
We then elaborate on the mentioned results by showing that if the $\omega$-categorical theory $T$ has “soft” elimination of imaginaries (Definition 2.9), which is a consequence of the more familiar concept of weak elimination of imaginaries [10, 9, 20, 6], and, for any $\mathcal{M}\models T$ and $C\subseteq M^{\mathrm{eq}}$, algebraic closure using parameters from $C$ restricted to elements of $M$ (i.e “real” elements) that have 0-rank 1 over $C$ has the exchange property, then $T$ is superrosy with finite $U^{\text{\rm\th }}$-rank; this is Theorem 6.14 below. It follows that if $T$ is $\omega$-categorical with weak elimination of imaginaries, $\mathcal{M}\models T$ and, for any $C\subseteq M^{\mathrm{eq}}$ the algebraic closure restricted to elements of $M$ that have 0-rank 1 over $C$ is trivial (Definition 6.11), then $T$ is superrosy with finite $U^{\text{\rm\th }}$-rank.
@@
Conant [6] has proved that the theory of the Fraïssé limit of any class of finite (relational) structures with free amalgamation has weak elimination of imaginaries and is superrosy with $U^{\text{\rm\th }}$-rank 1. The class of finite boolean algebras does not have the free amalgamation but it has the amalgamation property, so it has a Fraïssé limit, namely the (unique up to isomorphism) infinite countable atomless boolean algebra, with complete theory $T_{B}$ say. Since $T_{B}$ has elimination of quantifiers it is straightforward to see that the algebraic closure on the real elements of any model of $T_{B}$ is trivial. By work of Newelski and Wencel [20], $T_{B}$ has weak elimination of imaginaries and it follows from Theorem 6.14 below that $T_{B}$ is superrosy with finite $U^{\text{\rm\th }}$-rank, which may be a new result. In any case, it follows that Theorem 6.14 below applies to a larger class of theories than the class of theories considered in [6]. It may also be the case that Theorem 6.10 or Theorem 6.14 can be applied to $\omega$-categorical theories obtained by Hrushovski’s method of construction of a “generic structure” for a class of finite structures (see e.g. [8]), but I have not investigated this.
@@ 2. Preliminaries
(ii) $a\notin\mathrm{acl}^{m}(C)$ then $a\notin\mathrm{acl}^{n}(C)$.
Since each $\mathcal{M}_{n}$ is interpretable in $\mathcal{M}$ the following result follows from [9, Theorem 7.3.8]:
@@
The first part above follows from the well-known theorem of Engeler, Ryll-Nardzewski, and Svenonius [9, Theorem 7.3.1]. The second part is a consequence of the fact that, since $B$ is finite, there is $n\leq m<\omega$ such that $B\subseteq M_{m}$ and $Th(\mathcal{M}_{m})$ is $\omega$-categorical.
@@
We say that $T$ has soft elimination of imaginaries if for all $a\in M^{\mathrm{eq}}$ there is a finite sequence $\bar{b}$ of elements from $M$ such that $a\in\mathrm{acl}^{\mathrm{eq}}(\bar{b})$ and $\mathrm{rng}(\bar{b})\subseteq\mathrm{acl}^{\mathrm{eq}}(a)$, or equivalently, $\mathrm{acl}^{\mathrm{eq}}(a)=\mathrm{acl}^{\mathrm{eq}}(\bar{b})$.
The more commonly used notion of weak elimination of imaginaries (where, with the notation of the definition above, it is required that $a$ belongs to the definable closure of $\bar{b}$) implies soft elimination of imaginaries. Weak elimination of imaginaries has been studied, for example, by Hodges, Hodkinson, and Macpherson, in [10] where they demonstrated, among other things, that the complete theories of dense linear order, the random (or Rado) graph, and the Fraïssé limit of the set of finite $K_{n}$-free graphs have weak elimination of imaginaries. Conditions that apply to $\omega$-categorical theories and imply weak elimination of imaginaries are given by [10, Lemma 6.5 and Proposition 8.2]. Newelski and Wencel [20] have proved that the complete theory of an infinite boolean algebra with only finitely many atoms has weak elimination of imaginaries. The above mentioned result about $K_{n}$-free graphs has been generalized by Conant [6] to a result saying that the theory of every Fraïssé limit of a class of finite structures with free amalgamation has weak elimination of imaginaries.
The backbone of the theory that will be developed does not use the assumption that $T$ has soft elimination of imaginaries, but, as we will see, the “hierarchy” of ranks and independence relations that we will study collapses to the bottom level under the assumption of soft elimination of imaginaries.
@@ 3. $n$-rank
    For any ordinal $\alpha$, $\mathrm{rk}^{n}(A/B)\geq\alpha+1$ if there is $a\in\mathrm{acl}^{n}(A)\setminus\mathrm{acl}^{n}(B)$ such that and $\mathrm{rk}^{n}(A/\{a\}\cup B)\geq\alpha$.
@@
Let $T$ be the theory which expresses that $E$ is an equivalence relation with infinitely many equivalence classes all of which are infinite. Let $\mathcal{M}\models T$ and let $a\in M$. It is well-known (and easy to show) that $T$ is $\omega$-categorical with elimination of quantifiers. From this it easily follows that, for all $A\subseteq M$, $\mathrm{acl}^{0}(A)=A$. Therefore $\mathrm{rk}^{0}(a)=1$. Let $a]_{E}$ denote the equivalence class of $a$ with respect to $E$ as a element of $M^{\mathrm{eq}}$, so $a]_{E}\in M_{1}$. Then $\mathrm{rk}^{1}(a/\{a,[a]_{E}\})\geq 0$, and as $a\in\mathrm{acl}^{1}(a)\setminus\mathrm{acl}^{1}(\emptyset)$ we get $\mathrm{rk}^{1}(a/\{[a]_{E}\})\geq 1$. Since $a]_{E}\in\mathrm{acl}^{1}(a)\setminus\mathrm{acl}^{1}(\emptyset)$ wet $\mathrm{rk}^{1}(a)\geq 2$. As $T$ has elimination of quantifiers it follows that $E$ is the only $\emptyset$-definable equivalence relation on $M$. Hence $\mathrm{acl}^{1}(a)=\{a,[a]_{E}\}$ and consequently $\mathrm{rk}^{1}(a)=2$. Let $b\in M$ be such that $b\neq a$ and $b]_{E}=[a]_{E}$. By arguing similarly as above it follows that $\mathrm{rk}^{0}(a/b)=\mathrm{rk}^{0}(a)=1$ and $\mathrm{rk}^{1}(a/b)=1<2=\mathrm{rk}^{1}(a)$.
@@
such that, for all $i=0,\ldots,k$, $a^{\prime}_{i}\notin\mathrm{acl}^{n}(\{a^{\prime}_{1},\ldots,a^{\prime}_{i-1}\}\cup B)$. Then $\mathrm{acl}^{n}(\{a^{\prime}_{0},\ldots,a^{\prime}_{k}\}\cup B)\subseteq\mathrm{acl}^{n}(\{a_{1},\ldots,a_{k}\}\cup B)$ so, for all $i=k+1,\ldots,\alpha$, $a_{i}\notin\mathrm{acl}^{n}(\{a^{\prime}_{0},\ldots,a^{\prime}_{k},a_{k+1},\ldots,a_{i-1}\}\cup B)$ By part (i), the sequence $a^{\prime}_{0},\ldots,a^{\prime}_{k},a_{k+1},\ldots,a_{r}$ witnesses that $\mathrm{rk}(A/B)\geq\alpha+1$ which contradicts the assumption. $\square$
@@
Suppose that $T$ has soft elimination of imaginaries. Let $\alpha<\omega$ and $A,B\subseteq M^{\mathrm{eq}}$.
@@
Proof. (i) Suppose that $\mathrm{rk}^{n}(A/B)\geq\alpha$. By Lemma 3.5, there are $a_{1},\ldots,a_{\alpha}\in\mathrm{acl}^{n}(A)$ such that, for all $k=1,\ldots,\alpha$, $a_{k}\notin\mathrm{acl}^{n}(\{a_{1},\ldots,a_{k-1}\}\cup B)$. By induction on $\alpha$ we prove that there are $a^{\prime}_{1},\ldots,a^{\prime}_{\alpha}\in\mathrm{acl}^{0}(A)$ such that $a^{\prime}_{k}\notin\mathrm{acl}^{0}(\{a^{\prime}_{1},\ldots,a^{\prime}_{k-1}\}\cup B)$ and $a^{\prime}_{k}\in\mathrm{acl}^{0}(a_{k})$ for all $k=1,\ldots,\alpha$, and it follows from Lemma 3.5 that $\mathrm{rk}^{0}(A/B)\geq\alpha$. If $\alpha=0$ there is nothing to prove, so suppose that $\alpha>0$. By the induction hypothesis there are $a^{\prime}_{1},\ldots,a^{\prime}_{\alpha-1}\in\mathrm{acl}^{0}(A)$ such that $a^{\prime}_{k}\notin\mathrm{acl}^{0}(\{a^{\prime}_{1},\ldots,a^{\prime}_{k-1}\}\cup B)$ and $a^{\prime}_{k}\in\mathrm{acl}^{0}(a_{k})$ for all $k=1,\ldots,\alpha-1$. Since $T$ is assumed to have soft elimination of imaginaries, there is a finite sequence $\bar{c}$ of elements from $M_{0}=M$ such that $\mathrm{acl}^{\mathrm{eq}}(a_{\alpha})=\mathrm{acl}^{\mathrm{eq}}(\bar{c})$. Since $a_{\alpha}\notin\mathrm{acl}(\{a_{1},\ldots,a_{\alpha-1}\}\cup B)$ we have $\mathrm{rng}(\bar{c})\not\subseteq\mathrm{acl}(\{a_{1},\ldots,a_{\alpha-1}\}\cup B)$. As $a^{\prime}_{k}\in\mathrm{acl}^{0}(a_{k})$ for all $k=1,\ldots,\alpha-1$ it follows that $\mathrm{rng}(\bar{c})\not\subseteq\mathrm{acl}^{0}(\{a^{\prime}_{1},\ldots,a^{\prime}_{\alpha-1}\}\cup B)$. Thus we can choose some element from $\bar{c}$, which we call $a^{\prime}_{\alpha}$, such that $a^{\prime}_{\alpha}\notin\mathrm{acl}^{0}(\{a^{\prime}_{1},\ldots,a^{\prime}_{\alpha-1}\}\cup B)$. Now Lemma 3.5 implies that $\mathrm{rk}^{0}(A/B)\geq\alpha$.
(ii) Suppose that $A$ is finite. By Lemma 3.7, $\mathrm{rk}^{n}(A/B)=\alpha$ for some $\alpha<\omega$. By Lemma 3.4, $\mathrm{rk}^{n}(A/B)\geq\mathrm{rk}^{0}(A/B)$, and by part (i) $rk^{n}(A/B)\leq\mathrm{rk}^{0}(A/B)$. $\square$
@@
Finally we consider part (d). Let $\beta=\mathrm{rk}^{n}(B/C)$ and $\alpha=\mathrm{rk}^{n}(A/B)$, so $\alpha,\beta<\omega$. By Lemma 3.5, there are $b_{1},\ldots b_{\beta}\in\mathrm{acl}^{n}(B)$ and $a_{1},\ldots,a_{s}\in\mathrm{acl}^{n}(A)$ such that, for all $k=1,\ldots,\beta$, $b_{k}\notin\mathrm{acl}^{n}(\{b_{1},\ldots,b_{k-1}\}\cup C)$, and for all $l=1,\ldots,\alpha$, $a_{l}\notin\mathrm{acl}^{n}(\{a_{1},\ldots,a_{l-1}\}\cup B)$. Moreover, $\beta$ and $\alpha$ are maximal such that such sequences exist. By he same lemma we have $\mathrm{acl}^{n}(B)\subseteq\mathrm{acl}^{n}(\{b_{1},\ldots,b_{\beta}\}\cup C)$ and hence $a_{l}\notin\mathrm{acl}^{n}(\{b_{1},\ldots,b_{\beta},a_{1},\ldots,a_{l-1}\}\cup C)$, for all $l=1,\ldots,\alpha$. Since $B\subseteq A$ we have $\mathrm{acl}^{n}(B)\subseteq\mathrm{acl}^{n}(A)$ and the same lemma again implies that $\mathrm{rk}(A/C)\geq\beta+\alpha$. $\square$

### v2

@@ Abstract.
For an $\omega$-categorical theory $T$ and model $\mathcal{M}$ of $T$ we define a hierarchy of ranks, the $n$-ranks for $n<\omega$ which only care about imaginary elements “up to level nn”, where level $n$ contains every element of $M$ and every imaginary element that is an equivalence class of an $\emptyset$-definable equivalence relation on $n$-tuples of elements from $M$. Using the $n$-rank we define the notion of $n$-independence. For all $n<\omega$, the $n$-independence relation restricted to $M_{n}$ has all properties of an independence relation according to Kim and Pillay [16] with the possible exception of the symmetry property. We prove that, given any $n<\omega$, if $\mathcal{M}\models T$ and the algebraic closure in $\mathcal{M}^{\text{\rm eq}}$ restricted to imaginary elements “up to level nn” which have $n$-rank 1 (over some set of parameters) satisfies the exchange property, then $n$-independence is symmetric and hence an independence relation when restricted to $M_{n}$. Then we show that if $n$-independence is symmetric for all $n<\omega$, then $T$ is rosy. An application of this is that if $T$ has geometric elimination of imaginaries and the algebraic closure in $\mathcal{M}$ restricted to elements of $M$ of 0-rank 1 (over some set of parameters from $M^{\text{\rm eq}}$) satisfies the exchange property, then $T$ is superrosy with finite $U^{\text{\rm\th }}$-rank.
@@ 1. Introduction
A variety of notions of rank and independence have played an important role in model theory at least since Morley’s influential work on uncountably categorical theories [20] in the 1960ies. Such notions have been central for developing more or less general (meta) theories which divide complete first-order theories into various classes. Shelah [25] used them in his development of stability theory by which theories can be classified into $\omega$-stable, superstable, stable, or unstable. Later Shelah’s classification theory, and its notion of independence, was shown, by Kim and Pillay [16, 15], to make sense in a wider context and the classes of simple and supersimple first-order theories were introduced. Yet later, more general notions of independence, including thorn-independence, have been studied by Onshuus, Ealy and Adler and the classes of rosy and superrosy theories have been introduced [1, 7, 22]. The class of rosy theories is quite diverse and includes, for example, all simple theories (which exclude a linearly ordered universe) and all o-minimal theories (which assume a linearly ordered universe) [29]. It is also the largest class of theories for which there is an independence relation which satisfies certain basic and natural properties [7]. The work on NIP theories [26] also uses notions of rank and independence (and NIP theories and rosy theories partially overlap).
@@
In this study we consider (only) $\omega$-categorical theories $T$ and investigate a hierarchy of ranks, the $n$-ranks for $n<\omega$, where the $n$-rank is defined entirely in terms of the algebraic closure operator on $\mathcal{M}^{\mathrm{eq}}$ (where $\mathcal{M}$ is a model of the theory) restricted to the set $M_{n}$ containing every real element and every imaginary element that corresponds to an equivalence class of an $\emptyset$-definable equivalence relation on $M^{k}=\underbrace{M\times\ldots\times M}_{\text{$k$ times}}$ where $k\leq n$. With the $n$-rank we define a notion of $n$-independence, denoted $\smile$$|^{n}$. Without further assumptions than $\omega$-categoricity, $n$-independence restricted to $M_{n}$ has all the properties of an independence relation according to [16] (possibly) except for the symmetry property. The main technical contribution of this study is to isolate a property, parametrized by $n$, which implies (in fact, is equivalent to) that $n$-independence is symmetric. The property in question (Assumption 5.1) is the exchange property of algebraic closure restricted to elements of $M_{n}$ with $n$-rank 1 (over some $C\subseteq M^{\mathrm{eq}}$). More precisely, if, for any $C\subseteq M^{\mathrm{eq}}$, the algebraic closure using parameters from $C$ restricted to elements of $M_{n}$ with $n$-rank 1 over $C$ has the exchange property, then $n$-independence is symmetric, and hence it is an independence relation in the sense of [16] when restricted to $M_{n}$, which is stated by Theorem 5.15. Using this we can relatively easily show the statement of Theorem 6.9 that if, for all $n<\omega$, the algebraic closure restricted to elements of $M_{n}$ with $n$-rank 1 has the exchange property, then the theory is rosy, which is proved by showing that thorn-independence has local character (see [7, Theorem 3.7]). Thus, an $\omega$-categorical theory which is not rosy must, for some $n<\omega$, have a model $\mathcal{M}$ and $C\subseteq M^{\mathrm{eq}}$ such that algebraic closure using parameters from $C$ restricted to elements of $M_{n}$ with $n$-rank 1 over $C$ does not have the exchange property. This may be useful for finding a dividing line between $\omega$-categorical rosy theories (e.g. dense linear order and the Frïssé limit of a class of finite relational structures with free amalgamation [6]) and $\omega$-categorical non-rosy theories (e.g. the Fraïssé limit of finite boolean algebras and $T^{*}_{feq}$ in [1, Example 2.6]). A consequence of Theorem 6.9 is that if algebraic closure on $M^{\mathrm{eq}}$ is trivial in the sense of Definition 6.10, then the theory is rosy, as stated by Theorem 6.12.
We then consider the effects of assuming geometric elimination of imaginaries (Definition 2.9) in the context studied here, where geometric elimination of imaginaries is a consequence of the perhaps more familiar concept of weak elimination of imaginaries [11, 10, 21, 6]. It will be shown that if the $\omega$-categorical theory has geometric elimination of imaginaries, then the hierarchy of $n$-independence notions, for $n<\omega$, collapses to the bottom level, that is for all $n<\omega$, $A,B,C\subseteq M^{\mathrm{eq}}$, $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{n}$}}B$ if and only if $A\underset{C}{\raisebox{-2.0pt}[5.0pt][0.0pt]{$\smile$}\hskip-6.8pt\raisebox{3.0pt}[5.0pt][0.0pt]{$|^{0}$}}B$.
@@
We use this to show that if the $\omega$-categorical theory $T$ has geometric elimination of imaginaries and, for any $\mathcal{M}\models T$ and $C\subseteq M^{\mathrm{eq}}$, algebraic closure using parameters from $C$ restricted to elements of $M$ (i.e “real” elements) that have 0-rank 1 over $C$ has the exchange property, then $T$ is superrosy with finite $U^{\text{\rm\th }}$-rank; this is Theorem 6.14 below. Hence Theorems 6.9 and 6.14 are generalizations, in the context of $\omega$-categorical theories, of the result that if a theory has geometric elimination of imaginaries and algebraic closure on all real elements (not just those of 0-rank 1 over some set) has the exchange property, then the theory is superrosy with $U^{\text{\rm\th }}$-rank 1 (this is [7, Theorem 4.12] where the authors also contribute the result to Gagelman [9] and Adler).
@@
Conant [6] has proved that the theory of the Fraïssé limit of any class of finite (relational) structures with free amalgamation has weak elimination of imaginaries and is superrosy with $U^{\text{\rm\th }}$-rank 1. Such a theory has trivial algebraic closure on real elements and therefore the algebraic closure on real elements satisfies the exchange property. So unless every theory to which Theorem 6.14 applies can be constructed as the Fraïssé limit of a class of relational structures with free amalgamation, then Theorem 6.14 applies to a larger class of theories than [6]. It may also be the case that Theorem 6.9 or Theorem 6.14 can be applied to $\omega$-categorical theories obtained by Hrushovski’s method of construction of a “generic structure” for a class of finite structures (see e.g. [8]), but I have not investigated this.
@@ 2. Preliminaries
(ii) If $a\in M_{m}$ and $a\notin\mathrm{acl}^{m}(C)$ then $a\notin\mathrm{acl}^{n}(C)$.
Since each $\mathcal{M}_{n}$ is interpretable in $\mathcal{M}$ the following result follows from [10, Theorem 7.3.8]:
@@
The first part above follows from the well-known theorem of Engeler, Ryll-Nardzewski, and Svenonius [10, Theorem 7.3.1]. The second part is a consequence of the fact that, since $B$ is finite, there is $n\leq m<\omega$ such that $B\subseteq M_{m}$ and $Th(\mathcal{M}_{m})$ is $\omega$-categorical.
@@
We say that $T$ has geometric elimination of imaginaries if for all $a\in M^{\mathrm{eq}}$ there is a finite sequence $\bar{b}$ of elements from $M$ such that $a\in\mathrm{acl}^{\mathrm{eq}}(\bar{b})$ and $\mathrm{rng}(\bar{b})\subseteq\mathrm{acl}^{\mathrm{eq}}(a)$, or equivalently, $\mathrm{acl}^{\mathrm{eq}}(a)=\mathrm{acl}^{\mathrm{eq}}(\bar{b})$.
The more commonly used notion of weak elimination of imaginaries (where, with the notation of the definition above, it is required that $a$ belongs to the definable closure of $\bar{b}$) implies geometric elimination of imaginaries. Weak elimination of imaginaries has been studied, for example, by Hodges, Hodkinson, and Macpherson, in [11] where they demonstrated, among other things, that the complete theories of dense linear order, the random (or Rado) graph, and the Fraïssé limit of the set of finite $K_{n}$-free graphs have weak elimination of imaginaries. Conditions that apply to $\omega$-categorical theories and imply weak elimination of imaginaries are given by [11, Lemma 6.5 and Proposition 8.2]. Newelski and Wencel [21] have proved that the complete theory of an infinite boolean algebra with only finitely many atoms has weak elimination of imaginaries. The above mentioned result about $K_{n}$-free graphs has been generalized by Conant [6] to a result saying that the theory of every Fraïssé limit of a class of finite structures with free amalgamation has weak elimination of imaginaries.
The backbone of the theory that will be developed does not use the assumption that $T$ has geometric elimination of imaginaries, but, as we will see, the “hierarchy” of ranks and independence relations that we will study collapses to the bottom level under the assumption of geometric elimination of imaginaries.
@@ 3. $n$-rank
    For any ordinal $\alpha$, $\mathrm{rk}^{n}(A/B)\geq\alpha+1$ if there is $a\in\mathrm{acl}^{n}(A)\setminus\mathrm{acl}^{n}(B)$ such that $\mathrm{rk}^{n}(A/\{a\}\cup B)\geq\alpha$.
@@
Let $T$ be the theory which expresses that $E$ is an equivalence relation with infinitely many equivalence classes all of which are infinite. Let $\mathcal{M}\models T$ and let $a\in M$. It is well-known (and easy to show) that $T$ is $\omega$-categorical with elimination of quantifiers. From this it easily follows that, for all $A\subseteq M$, $\mathrm{acl}^{0}(A)=A$. Therefore $\mathrm{rk}^{0}(a)=1$. Let $a]_{E}$ denote the equivalence class of $a$ with respect to $E$ as a element of $M^{\mathrm{eq}}$, so $a]_{E}\in M_{1}$. Then $\mathrm{rk}^{1}(a/\{a,[a]_{E}\})\geq 0$, and as $a\in\mathrm{acl}^{1}(a)\setminus\mathrm{acl}^{1}(\emptyset)$ we get $\mathrm{rk}^{1}(a/\{[a]_{E}\})\geq 1$. Since $a]_{E}\in\mathrm{acl}^{1}(a)\setminus\mathrm{acl}^{1}(\emptyset)$ we get $\mathrm{rk}^{1}(a)\geq 2$. As $T$ has elimination of quantifiers it follows that $E$ is the only $\emptyset$-definable equivalence relation on $M$. Hence $\mathrm{acl}^{1}(a)=\{a,[a]_{E}\}$ and consequently $\mathrm{rk}^{1}(a)=2$. Let $b\in M$ be such that $b\neq a$ and $b]_{E}=[a]_{E}$. By arguing similarly as above it follows that $\mathrm{rk}^{0}(a/b)=\mathrm{rk}^{0}(a)=1$ and $\mathrm{rk}^{1}(a/b)=1<2=\mathrm{rk}^{1}(a)$.
@@
such that, for all $i=0,\ldots,k$, $a^{\prime}_{i}\notin\mathrm{acl}^{n}(\{a^{\prime}_{1},\ldots,a^{\prime}_{i-1}\}\cup B)$. Then $\mathrm{acl}^{n}(\{a^{\prime}_{0},\ldots,a^{\prime}_{k}\}\cup B)\subseteq\mathrm{acl}^{n}(\{a_{1},\ldots,a_{k}\}\cup B)$ so, for all $i=k+1,\ldots,\alpha$, $a_{i}\notin\mathrm{acl}^{n}(\{a^{\prime}_{0},\ldots,a^{\prime}_{k},a_{k+1},\ldots,a_{i-1}\}\cup B)$ By part (i), the sequence $a^{\prime}_{0},\ldots,a^{\prime}_{k},a_{k+1},\ldots,a_{\alpha}$ witnesses that $\mathrm{rk}(A/B)\geq\alpha+1$ which contradicts the assumption. $\square$
@@
Suppose that $T$ has geometric elimination of imaginaries. Let $\alpha<\omega$ and $A,B\subseteq M^{\mathrm{eq}}$.
@@
Proof. (i) Suppose that $\mathrm{rk}^{n}(A/B)\geq\alpha$. By Lemma 3.5, there are $a_{1},\ldots,a_{\alpha}\in\mathrm{acl}^{n}(A)$ such that, for all $k=1,\ldots,\alpha$, $a_{k}\notin\mathrm{acl}^{n}(\{a_{1},\ldots,a_{k-1}\}\cup B)$. By induction on $\alpha$ we prove that there are $a^{\prime}_{1},\ldots,a^{\prime}_{\alpha}\in\mathrm{acl}^{0}(A)$ such that $a^{\prime}_{k}\notin\mathrm{acl}^{0}(\{a^{\prime}_{1},\ldots,a^{\prime}_{k-1}\}\cup B)$ and $a^{\prime}_{k}\in\mathrm{acl}^{0}(a_{k})$ for all $k=1,\ldots,\alpha$, and it follows from Lemma 3.5 that $\mathrm{rk}^{0}(A/B)\geq\alpha$. If $\alpha=0$ there is nothing to prove, so suppose that $\alpha>0$. By the induction hypothesis there are $a^{\prime}_{1},\ldots,a^{\prime}_{\alpha-1}\in\mathrm{acl}^{0}(A)$ such that $a^{\prime}_{k}\notin\mathrm{acl}^{0}(\{a^{\prime}_{1},\ldots,a^{\prime}_{k-1}\}\cup B)$ and $a^{\prime}_{k}\in\mathrm{acl}^{0}(a_{k})$ for all $k=1,\ldots,\alpha-1$. Since $T$ is assumed to have geometric elimination of imaginaries, there is a finite sequence $\bar{c}$ of elements from $M_{0}=M$ such that $\mathrm{acl}^{\mathrm{eq}}(a_{\alpha})=\mathrm{acl}^{\mathrm{eq}}(\bar{c})$. Since $a_{\alpha}\notin\mathrm{acl}^{n}(\{a_{1},\ldots,a_{\alpha-1}\}\cup B)$ we have $\mathrm{rng}(\bar{c})\not\subseteq\mathrm{acl}^{n}(\{a_{1},\ldots,a_{\alpha-1}\}\cup B)$. As $a^{\prime}_{k}\in\mathrm{acl}^{0}(a_{k})$ for all $k=1,\ldots,\alpha-1$ it follows that $\mathrm{rng}(\bar{c})\not\subseteq\mathrm{acl}^{0}(\{a^{\prime}_{1},\ldots,a^{\prime}_{\alpha-1}\}\cup B)$. Thus we can choose some element from $\bar{c}$, which we call $a^{\prime}_{\alpha}$, such that $a^{\prime}_{\alpha}\notin\mathrm{acl}^{0}(\{a^{\prime}_{1},\ldots,a^{\prime}_{\alpha-1}\}\cup B)$. Now Lemma 3.5 implies that $\mathrm{rk}^{0}(A/B)\geq\alpha$.
(ii) Suppose that $A$ is finite. By Lemma 3.7, $\mathrm{rk}^{n}(A/B)=\alpha$ for some $\alpha<\omega$. By Lemma 3.4, $\mathrm{rk}^{n}(A/B)\geq\mathrm{rk}^{0}(A/B)$, and by part (i) $\mathrm{rk}^{n}(A/B)\leq\mathrm{rk}^{0}(A/B)$. $\square$
@@
Finally we consider part (d). Let $\beta=\mathrm{rk}^{n}(B/C)$ and $\alpha=\mathrm{rk}^{n}(A/BC)$, so $\alpha,\beta<\omega$. By Lemma 3.5, there are $b_{1},\ldots b_{\beta}\in\mathrm{acl}^{n}(B)$ and $a_{1},\ldots,a_{\alpha}\in\mathrm{acl}^{n}(A)$ such that, for all $k=1,\ldots,\beta$, $b_{k}\notin\mathrm{acl}^{n}(\{b_{1},\ldots,b_{k-1}\}\cup C)$, and for all $l=1,\ldots,\alpha$, $a_{l}\notin\mathrm{acl}^{n}(\{a_{1},\ldots,a_{l-1}\}\cup BC)$. Moreover, $\beta$ and $\alpha$ are maximal such that such sequences exist. For a contradiction, suppose that, for some $l\in\{1,\ldots,\alpha\}$, $a_{l}\in\mathrm{acl}^{n}(\{b_{1},\ldots,b_{\beta},a_{1},\ldots,a_{l-1}\}\cup C)$. Since $b_{1},\ldots,b_{\beta}\in\mathrm{acl}^{n}(B)$ we get $a_{l}\in\mathrm{acl}^{n}(\{a_{1},\ldots,a_{l-1}\}\cup BC)$ which contradicts the choice of $a_{1},\ldots,a_{\alpha}$. Hence we conclude that $a_{l}\notin\mathrm{acl}^{n}(\{b_{1},\ldots,b_{\beta},a_{1},\ldots,a_{l-1}\}\cup C)$ for all $l=1,\ldots,\alpha$. Since $B\subseteq A$ we have $\mathrm{acl}^{n}(B)\subseteq\mathrm{acl}^{n}(A)$ and therefore $b_{1},\ldots,b_{\beta},a_{1},\ldots,a_{\alpha}\in\mathrm{acl}^{n}(A)$. By Lemma 3.5, we get $\mathrm{rk}(A/C)\geq\beta+\alpha$. $\square$